# From Clicks to Preference: A Multi-stage Alignment Framework for Generative Query Suggestion in Conversational System

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: KDD; August 09–13, 2026; Jeju, Koreaisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Query suggestionccs: Computing methodologies Natural language generation
@@ 1. Introduction
The proliferation of conversational systems, including chatbots and virtual assistants, has significantly reshaped the landscape of human-computer interaction (Gao et al., 2018). These systems are rapidly moving beyond niche applications to become ubiquitous platforms for information access, task completion, and customer support (Shum et al., 2018). With this evolving landscape, query suggestion (QS) has emerged as a cornerstone of effective conversational user experience. As illustrated in Figure 1, its primary function is to proactively recommend relevant queries based on the conversational context, making the interaction faster and more satisfying. Ideally, an intelligent QS module should not merely autocomplete user queries, but actively lead the conversation toward a success outcome.
@@
Traditional methods in QS, including collaborative filtering (Schafer et al., 2007) and session-based recommendations (Chen and Chen, 2021; Ludewig et al., 2021; Ludewig and Jannach, 2018), operate on a retrieval-based paradigm, selecting suggestions from a finite candidate pool of previously observed queries. This design makes them inherently unable to generate novel, contextually specific suggestions and leaves them vulnerable to chronic issues like data sparsity and the cold-start problem (Fayyaz et al., 2020). The advent of Large Language Models (LLMs) (Achiam et al., 2023; Team et al., 2025; Guo et al., 2025) has catalyzed a paradigm shift from retrieval-based matching to semantic generation, directly addressing these foundational constraints with their superior contextual understanding and world knowledge (Wang et al., 2023; He et al., 2023; Wang et al., 2024a).
@@
In the QS context, an additional challenge is that user choice is non-deterministic and noisy, making a reward model based on a deterministic score imprecise. To this end, we propose the Gaussian Reward Model (GaRM), which conceptualizes preference scores not as deterministic scalars but as probabilistic distributions. Critically, we derive an analytically tractable loss function with variance regularization for GaRM, overcoming the inefficiency and instability of prior probabilistic approaches (Sun et al., 2025) and making it robust to the inherent noise in preference data.
@@ 2.1. Query Suggestion in Search Engine
Query Suggestion (QS) has been extensively studied within the context of traditional search engine, where its primary goal is to assist users by recommending queries based on historical usage patterns. Early QS research relied on mining query co-occurrence in logs, click-through statistics, and simple probabilistic models such as n-grams or Markov chains (Baeza-Yates et al., 2004; Cao et al., 2008). While effective for popular or repetitive queries, such approaches fall short in open-domain conversations due to limited semantic understanding and inability to handle unseen inputs. To overcome these limitations, researches introduced neural network-based methods, particularly recurrent neural networks and attention-based models (Sordoni et al., 2015; Mustar et al., 2020, 2021; Dehghani et al., 2017), to better capture sequential dependencies in session data and generalize across diverse queries. However, these systems are still constrained by a fixed set of candidate queries and require extensive offline training data.
@@ 2.2. Query Suggestion in Conversational Systems
In contrast to traditional search QS, conversational QS aims to proactively guide the conversation toward more fruitful, engaging, or task-completing directions, a concept termed ”conversation-leading” suggestions (Rosset et al., 2020). This capability has become increasingly central in large-scale conversational AI systems like ChatGPT (Ouyang et al., 2022), Claude (Bai et al., 2022), and Doubao (Seed, 2025), where users often rely on system-initiated suggestions to navigate complex, multi-turn interactions. Rosset et al. (Rosset et al., 2020) pioneered the notion of conversation-leading query suggestions—informative prompts that anticipate user needs rather than merely continuing the current thought. Wang et al. (Wang et al., 2023) further demonstrated the efficacy of zero-shot LLMs in generating high-quality query continuations. Such generative approaches are particularly valuable in cold-start or exploratory settings, but they also introduce challenges. Without proper alignment, LLMs may hallucinate content, repeat generic suggestions, or deviate from useful dialogue trajectories.
@@ 2.3. Human Preference Alignment
Aligning generative models with nuanced human preferences is a central challenge in modern AI. The dominant paradigm for this is Reinforcement Learning from Human Feedback (RLHF), which consists of three main stages: SFT, reward model (RM) training, and RL-based policy optimization. The core idea of learning from human preferences was pioneered by Christiano et al. (Christiano et al., 2017), and the full RLHF pipeline was popularized by its successful application to large-scale language models like InstructGPT (Ouyang et al., 2022). In the context of QS, the primary challenge is defining a suitable preference signal. Consequently, research has focused on using implicit user feedback, such as clicks, as a proxy for preference (Min et al., 2025a). For instance, Min et al. designed a system that first train a click-through rate (CTR) predictor to act as a reward model and then use it to guide the generative model toward suggestions that are more likely to be clicked (Min et al., 2025a). Furthermore, the authors used the predicted CTR to weight the importance of preference pairs in the Direct Preference Optimization (DPO) loss function and combined it with a diversity-aware regularization term to ensure that the suggestions remain varied (Min et al., 2025b). However, relying on CTR models as preference proxies presents significant challenges. Primarily, such models tend to associate clicks with the input context, rather than modeling the relative quality differences between candidate suggestions. Furthermore, user clicks are stochastic by nature, but a simple CTR model cannot represent this probabilistic uncertainty, treating each signal as a deterministic event. In response to these issues, we propose a probabilistic, pairwise GaRM to explicitly model the uncertainty in user preference. We then employ a full RL stage, complete with a composite reward and novel regularization, to achieve a more robust and holistic alignment with true user satisfaction.
@@ 3.4. Supervised Fine-tuning
High-Quality Candidate Synthesis. We begin by training a powerful, large-scale ”teacher” model (e.g., Gemma3-27B (Team et al., 2025), Qwen3-32B (Team, 2025)) on a cleaned, single-suggestion dataset. Freed from online inference constraints, this model is optimized solely for generation quality. We then use this teacher model to perform parallel sampling for each source dialogue context, generating a large candidate pool of high-quality individual suggestions.
@@ 3.5. Preference Modeling
Bradley-Terry Reward Model. This is a classical modeling approach that maps each suggestion and its context into a scalar score $r_{\phi}(s;h)$, which we use as a baseline in this work (Ouyang et al., 2022; Wang et al., 2024b). The loss function is defined as:
@@
This approach is known as PairedRM in previous literature (Park et al., 2024; Liu et al., 2025; Jiang et al., 2023). Further implementation details are provided in Appendix E.
Gaussian Reward Model. Due to the inherent variability in user preferences, the dataset of suggestion preferences contains significant noise that is difficult to eliminate through filtering alone. Therefore, it is crucial to design a more robust reward model at the algorithmic level. To this end, we propose modeling user preference for a suggestion as a probability distribution rather than a scalar value, as in BTRM. Inspired by and improving upon PURM (Sun et al., 2025), we model preferences as a parameterized distribution $\mathcal{N}(\mu,\sigma^{2})$. In the original PURM formulation, the optimization objective is to maximize:
@@ 3.6. Reinforcement Learning
To address these challenges, we propose leveraging multiple reward signals and jointly optimizing them during reinforcement learning. We categorize these reward signals into four types. Rule-based reward refers to programmatically defined scoring functions applied to generated texts, including metrics such as diversity checks and language consistency. Prompt-based reward uses API calls to LLMs, where we provide instructions defining the scoring criteria and let the LLM assign a score. Regularization reward is introduced to discourage the fine-tuned model from deviating excessively from the original model; in this work, we use the logged perplexity (PPL) as a regularization term. Neural network-based reward includes the three types of reward models trained in the previous section, along with the pretrained reward model provided by Skywork (Liu et al., 2024). We briefly discuss several types below, and more information can be found in Appendix H.
@@
Aligning with recent findings, which demonstrate that rule-based ’checklists’ are an effective strategy for reward engineering (Viswanathan et al., 2025), we employ several verifiable rewards to ensure more controllable and desirable model behavior. These include:
@@
Length Reward. To encourage conciseness, we penalize suggestions that are excessively long. We set a soft limit for each suggestion at a length of 12 words. If a suggestion surpasses this limit, a penalty is applied that scales linearly with the number of excess words (Yu et al., 2025).
@@ 3.6.4. RL Optimization
We adopt Group Relative Policy Optimization (GRPO) (Shao et al., 2024) to maximize the reward signal. GRPO optimizes the policy $\pi_{\theta}$ by maximizing a penalized objective that balances reward maximization against policy divergence. For each sample $i$ with input $h_{i}$, the policy generates an output $s_{i}$. The update leverages a reward signal $R(h_{i},s_{i})$ and constrains the policy shift using the Kullback-Leibler divergence. The objective is formulated as:
@@ 4.1. Setup
Our framework is implemented and evaluated within a large-scale, production conversational AI system–Cici. To ensure the models are trained and tested on representative data, all datasets are sampled from anonymized, real-world user interactions with the live system. We adopt the Qwen3-30B-A3B foundation model (Team, 2025) for its state-of-the-art performance and computational efficiency—two key requirements for deployment in a production environment.
@@ 4.2. Metrics
We assess the quality of each generated suggestion group (containing up to three items) by assigning an integer score from 0 to 3, corresponding to the number of useful suggestions. Scoring is conducted through a hybrid process: initial annotations are performed by gpt-o3-mini (OpenAI, [n. d.]), followed by a manual review by six human experts. Our offline metric, Good Same Bad (GSB), is then calculated as the aggregate score difference between a candidate model and a baseline across the test set.
@@
To assess the real-world performance of our proposed strategies, we conduct online A/B testing over a seven-day period on our conversational AI platform, Cici. In this experiment, each strategy is deployed as a distinct variant and is allocated 5% of the total user traffic for evaluation. The primary performance metric is CTR, defined as the total number of clicks on suggestion groups divided by their total number of impressions.
@@ 4.3.2. Safety Evaluation
The SFT stage yields a substantial initial improvement, elevating safety accuracy from 75.5% to 88.0%. Subsequently, all RL strategies, built on SFT model and guided by a format-based reward that penalizes improper outputs, further improve this metric to over 90%.
@@ 4.3.3. Reward Model Evaluation
Furthermore, we calculate the ULB for each individual sample pair within these datasets. We then group the samples into bins based on their confidence scores and compute the prediction accuracy of GaRM for each bin. As illustrated in Figure 4, there is a clear positive trend: the higher the confidence score of a sample, the higher the model’s accuracy on that sample. This demonstrates that the confidence score learned by GaRM is not arbitrary but carries a tangible and meaningful interpretation of the model’s certainty.
@@ 5. Conclusion
In this work, we propose a multi-stage framework for conversational query suggestion, which integrates prompt engineering, supervised fine-tuning, user preference modeling, and reinforcement learning. The proposed framework progressively aligns the generation policy to real user preference. Our framework achieves a relative click-through rate improvement of over 30% and, more importantly, establishes a virtuous cycle where improved models generate higher-quality data for future training iterations. In future, we will focus on enhancing the reward model by constructing targeted preference datasets to boost its in-distribution accuracy and by exploring new techniques to strengthen its out-of-distribution generalization.
@@ Appendix E Detailed Implementation for PairedRM
To better leverage the characteristics of autoregressive LLMs in constructing PairedRM, we adopt the method from Qwen3-embedding  (Zhang et al., 2025) by incorporating an instruction template into the input data. This template includes a task description, historical context, and the two suggestions to be evaluated. The specific template is shown in Figure 5.
@@ H.2. PE-based Reward
Acknowledging that having an LLM directly score nuanced user preferences is unreliable, we adopt a more robust, rubric-based approach. We prompt a powerful LLM with a set of clear, verifiable criteria (e.g., stylistic requirements, structural constraints) and task it with returning a binary quality score (1 for high, 0 for low). This strategy of using an LLM as a proxy evaluator has proven effective in prior work (Gunjal et al., 2025) and provides a scalable reward signal for our reinforcement learning loop. The prompt template is shown in Figure 7.

### v2

@@ Abstract.
journalyear: 2026copyright: ccconference: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 09–13, 2026; Jeju Island, Koreabooktitle: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’26), August 09–13, 2026, Jeju Island, Republic of Koreadoi: 10.1145/3770854.3783953isbn: 979-8-4007-2258-5/2026/08ccs: Information systems Query suggestionccs: Computing methodologies Natural language generation
@@ 1. Introduction
The proliferation of conversational systems, including chatbots and virtual assistants, has significantly reshaped the landscape of human-computer interaction (gao2018neural). These systems are rapidly moving beyond niche applications to become ubiquitous platforms for information access, task completion, and customer support (shum2018eliza). With this evolving landscape, query suggestion (QS) has emerged as a cornerstone of effective conversational user experience. As illustrated in Figure 1, its primary function is to proactively recommend relevant queries based on the conversational context, making the interaction faster and more satisfying. Ideally, an intelligent QS module should not merely autocomplete user queries, but actively lead the conversation toward a success outcome.
@@
Traditional methods in QS, including collaborative filtering (schafer2007collaborative) and session-based recommendations (chen2021collaborative; ludewig2021empirical; ludewig2018evaluation), operate on a retrieval-based paradigm, selecting suggestions from a finite candidate pool of previously observed queries. This design makes them inherently unable to generate novel, contextually specific suggestions and leaves them vulnerable to chronic issues like data sparsity and the cold-start problem (fayyaz2020recommendation). The advent of Large Language Models (LLMs) (achiam2023gpt; gemma_2025; guo2025deepseek) has catalyzed a paradigm shift from retrieval-based matching to semantic generation, directly addressing these foundational constraints with their superior contextual understanding and world knowledge (wang2023zero; he2023large; wang2024multimodal).
@@
In the QS context, an additional challenge is that user choice is non-deterministic and noisy, making a reward model based on a deterministic score imprecise. To this end, we propose the Gaussian Reward Model (GaRM), which conceptualizes preference scores not as deterministic scalars but as probabilistic distributions. Critically, we derive an analytically tractable loss function with variance regularization for GaRM, overcoming the inefficiency and instability of prior probabilistic approaches (sun2025probabilistic) and making it robust to the inherent noise in preference data.
@@ 2.1. Query Suggestion in Search Engine
Query Suggestion (QS) has been extensively studied within the context of traditional search engine, where its primary goal is to assist users by recommending queries based on historical usage patterns. Early QS research relied on mining query co-occurrence in logs, click-through statistics, and simple probabilistic models such as n-grams or Markov chains (baeza2004query; cao2008context). While effective for popular or repetitive queries, such approaches fall short in open-domain conversations due to limited semantic understanding and inability to handle unseen inputs. To overcome these limitations, researches introduced neural network-based methods, particularly recurrent neural networks and attention-based models (sordoni2015hierarchical; mustar2020using; mustar2021study; dehghani2017learning), to better capture sequential dependencies in session data and generalize across diverse queries. However, these systems are still constrained by a fixed set of candidate queries and require extensive offline training data.
@@ 2.2. Query Suggestion in Conversational Systems
In contrast to traditional search QS, conversational QS aims to proactively guide the conversation toward more fruitful, engaging, or task-completing directions, a concept termed ”conversation-leading” suggestions (rosset2020leading). This capability has become increasingly central in large-scale conversational AI systems like ChatGPT (ouyang2022training), Claude (bai2022constitutional), and Doubao (seed2025th), where users often rely on system-initiated suggestions to navigate complex, multi-turn interactions. Rosset et al. (rosset2020leading) pioneered the notion of conversation-leading query suggestions—informative prompts that anticipate user needs rather than merely continuing the current thought. Wang et al. (wang2023zero) further demonstrated the efficacy of zero-shot LLMs in generating high-quality query continuations and Li et al. (Li2025ProactiveGO) train a language model to enhance search chat-bot. Such generative approaches are particularly valuable in cold-start or exploratory settings, but they also introduce challenges. Without proper alignment, LLMs may hallucinate content, repeat generic suggestions, or deviate from useful dialogue trajectories.
@@ 2.3. Human Preference Alignment
Aligning generative models with nuanced human preferences is a central challenge in modern AI. The dominant paradigm for this is Reinforcement Learning from Human Feedback (RLHF), which consists of three main stages: SFT, reward model (RM) training, and RL-based policy optimization. The core idea of learning from human preferences was pioneered by Christiano et al. (christiano2017deep), and the full RLHF pipeline was popularized by its successful application to large-scale language models like InstructGPT (ouyang2022training). In the context of QS, the primary challenge is defining a suitable preference signal. Consequently, research has focused on using implicit user feedback, such as clicks, as a proxy for preference (min2025prompting). For instance, Min et al. designed a system that first train a click-through rate (CTR) predictor to act as a reward model and then use it to guide the generative model toward suggestions that are more likely to be clicked (min2025prompting). Furthermore, the authors used the predicted CTR to weight the importance of preference pairs in the Direct Preference Optimization (DPO) loss function and combined it with a diversity-aware regularization term to ensure that the suggestions remain varied (min2025ctr). However, relying on CTR models as preference proxies presents significant challenges. Primarily, such models tend to associate clicks with the input context, rather than modeling the relative quality differences between candidate suggestions. Furthermore, user clicks are stochastic by nature, but a simple CTR model cannot represent this probabilistic uncertainty, treating each signal as a deterministic event. In response to these issues, we propose a probabilistic, pairwise GaRM to explicitly model the uncertainty in user preference. We then employ a full RL stage, complete with a composite reward and novel regularization, to achieve a more robust and holistic alignment with true user satisfaction.
@@ 3.4. Supervised Fine-tuning
High-Quality Candidate Synthesis. We begin by training a powerful, large-scale ”teacher” model (e.g., Gemma3-27B (gemma_2025), Qwen3-32B (qwen3technicalreport)) on a cleaned, single-suggestion dataset. Freed from online inference constraints, this model is optimized solely for generation quality. We then use this teacher model to perform parallel sampling for each source dialogue context, generating a large candidate pool of high-quality individual suggestions.
@@ 3.5. Preference Modeling
Bradley-Terry Reward Model. This is a classical modeling approach that maps each suggestion and its context into a scalar score $r_{\phi}(s;h)$, which we use as a baseline in this work (ouyang2022training; wang2024secrets). The loss function is defined as:
@@
This approach is known as PairedRM in previous literature (pairedrm1; pairedrm2; pairdrm3). Further implementation details are provided in Appendix E.
Gaussian Reward Model. Due to the inherent variability in user preferences, the dataset of suggestion preferences contains significant noise that is difficult to eliminate through filtering alone. Therefore, it is crucial to design a more robust reward model at the algorithmic level. To this end, we propose modeling user preference for a suggestion as a probability distribution rather than a scalar value, as in BTRM. Inspired by and improving upon PURM (sun2025probabilistic), we model preferences as a parameterized distribution $\mathcal{N}(\mu,\sigma^{2})$. In the original PURM formulation, the optimization objective is to maximize:
@@ 3.6. Reinforcement Learning
To address these challenges, we propose leveraging multiple reward signals and jointly optimizing them during reinforcement learning. We categorize these reward signals into four types. Rule-based reward refers to programmatically defined scoring functions applied to generated texts, including metrics such as diversity checks and language consistency. Prompt-based reward uses API calls to LLMs, where we provide instructions defining the scoring criteria and let the LLM assign a score. Regularization reward is introduced to discourage the fine-tuned model from deviating excessively from the original model; in this work, we use the logged perplexity (PPL) as a regularization term. Neural network-based reward includes the three types of reward models trained in the previous section, along with the pretrained reward model provided by Skywork (liu2024skywork). We briefly discuss several types below, and more information can be found in Appendix H.
@@
Aligning with recent findings, which demonstrate that rule-based ’checklists’ are an effective strategy for reward engineering (viswanathan2025checklists), we employ several verifiable rewards to ensure more controllable and desirable model behavior. These include:
@@
Length Reward. To encourage conciseness, we penalize suggestions that are excessively long. We set a soft limit for each suggestion at a length of 12 words. If a suggestion surpasses this limit, a penalty is applied that scales linearly with the number of excess words (yu2025dapo).
@@ 3.6.4. RL Optimization
We adopt Group Relative Policy Optimization (GRPO) (shao2024deepseekmath) to maximize the reward signal. GRPO optimizes the policy $\pi_{\theta}$ by maximizing a penalized objective that balances reward maximization against policy divergence. For each sample $i$ with input $h_{i}$, the policy generates an output $s_{i}$. The update leverages a reward signal $R(h_{i},s_{i})$ and constrains the policy shift using the Kullback-Leibler divergence. The objective is formulated as:
@@ 4.1. Setup
Our framework is implemented and evaluated within a large-scale, production conversational AI system–Cici. To ensure the models are trained and tested on representative data, all datasets are sampled from anonymized, real-world user interactions with the live system. We adopt the Qwen3-30B-A3B foundation model (qwen3technicalreport) for its state-of-the-art performance and computational efficiency—two key requirements for deployment in a production environment.
@@ 4.2. Metrics
We assess the quality of each generated suggestion group (containing up to three items) by assigning an integer score from 0 to 3, corresponding to the number of useful suggestions. Scoring is conducted through a hybrid process: initial annotations are performed by gpt-o3-mini (openai_o3_mini_system_card), followed by a manual review by six human experts. Our offline metric, Good Same Bad (GSB), is then calculated as the aggregate score difference between a candidate model and a baseline across the test set.
@@
To assess the real-world performance of our proposed strategies, we conduct online A/B testing over a seven-day period on our conversational AI platform, Cici. In this experiment, each strategy is deployed as a distinct variant and is allocated 5% of the total user traffic for evaluation, involving millions of users per arm. The primary performance metric is CTR, defined as the total number of clicks on suggestion groups divided by their total number of impressions. All experiments yield statistically significant CTR gains (p¡0.05).
@@ 4.3.2. Safety Evaluation
The SFT stage yields a substantial initial improvement, elevating safety accuracy from 75.5% to 88.0%. Subsequently, all RL strategies, built on SFT model and guided by a format-based reward that penalizes improper outputs, further improve this metric score to over 90%.
@@ 4.3.3. Reward Model Evaluation
Furthermore, we calculate the ULB for each individual sample pair within these datasets. We then group the samples into bins based on their confidence scores and compute the prediction accuracy of GaRM for each bin. As illustrated in Figure 4, there is a clear positive trend: the higher the confidence score of a sample, the higher the model’s accuracy on that sample. This demonstrates that the confidence score learned by GaRM is not arbitrary but carries a tangible and meaningful interpretation of the model’s certainty. We also calculate ECE (Expected Calibration Error (Guo2017OnCO)) for a calibration assessment. For the Week 4 dataset, the ECE is 0.0302, while for the Week 4-RFT dataset, it is 0.0886. Both scores are low, with the larger ECE observed on the dataset that exhibits a stronger domain shift. This validates that ULB score can precisely quantify the uncertainty level.
4.3.4. Deployment.
We adopt 8-bit quantization and deploy the model using an in-house LLM Server. Each model copy is deployed on an individual GPU (only data parallel), achieving an average response time of 800ms. This speed exceeds the user’s reading speed, making it entirely feasible for a good user experience. The total number of GPUs supporting our online service is flexible according to online traffic.
@@ 4.4. Ablation Study
DPO-GaRM +69
We also compare a Direct Preference Optimization (DPO) strategy (Rafailov2023DirectPO), which uses all reward functions to construct positive and negative samples for training. The results show that the GSB score drops by 11 points, validating the superiority of the GRPO component.
@@ 5. Conclusion
In this work, we propose a multi-stage framework for conversational query suggestion, which integrates prompt engineering, supervised fine-tuning, user preference modeling, and reinforcement learning. The proposed framework progressively aligns the generation policy to real user preference. Our framework achieves a relative click-through rate improvement of over 30% and, more importantly, establishes a virtuous cycle where improved models generate higher-quality data for future training iterations. Notably, our method also indirectly improved key secondary metrics, such as the number of user messages (+1.2%) and active days per user (+1.16%), both with statistical significance. In future, we will focus on enhancing the reward model by constructing targeted preference datasets to boost its in-distribution accuracy and by exploring new techniques to strengthen its out-of-distribution generalization.
@@ Appendix E Detailed Implementation for PairedRM
To better leverage the characteristics of autoregressive LLMs in constructing PairedRM, we adopt the method from Qwen3-embedding  (zhang2025qwen3) by incorporating an instruction template into the input data. This template includes a task description, historical context, and the two suggestions to be evaluated. The specific template is shown in Figure 5.
@@ H.2. PE-based Reward
Acknowledging that having an LLM directly score nuanced user preferences is unreliable, we adopt a more robust, rubric-based approach. We prompt a powerful LLM with a set of clear, verifiable criteria (e.g., stylistic requirements, structural constraints) and task it with returning a binary quality score (1 for high, 0 for low). This strategy of using an LLM as a proxy evaluator has proven effective in prior work (gunjal2025rubrics) and provides a scalable reward signal for our reinforcement learning loop. The prompt template is shown in Figure 7.